# Exploring Neural Granger Causality with xLSTMs: Unveiling Temporal Dependencies in Complex Data

### v4

@@ 5 Conclusion
This work received funding from the ACATIS Investment KVG mbH project “Temporal Machine Learning for Long-Term Value Investing” and the KompAKI project of the Federal Ministry of Education and Research of Germany within the “The Future of Value Creation – Research on Production, Services and Work” program (funding number 02L19C150), managed by the Project Management Agency Karlsruhe (PTKA). The author of Eindhoven University of Technology received support from their Department of Mathematics and Computer Science and the Eindhoven Artificial Intelligence Systems Institute. Furthermore, this work benefited from the DYNAMIC Centre funded by the LOEWE program of the Hessian Ministry of Science and Research, Arts and Culture (HMWK) as LOEWE 1/16/519/03/09.001 (0009)98 and the project “The Third Wave of Artificial Intelligence - 3AI” of the HMWK. Map data © OpenStreetMap contributors, licensed under the ODbL and available from openstreetmap.org.
@@ Appendix A Full Algorithm
Algorithm 1 describes the entire procedure for extracting a Granger causal graph using GC-xLSTM. $(\cdot)_{+}$ denotes truncation as $\max(\cdot,0)$.
@@
5:     $\bm{\phi}_{v}:\bm{W}_{v},\bm{b}_{v}\sim\mathcal{U}\left(-\sqrt{\nicefrac{{1}}{{V}}},\sqrt{\nicefrac{{1}}{{V}}}\right)$triangleright Kaiming/He initialization
6:     $\bm{\alpha}_{v}\leftarrow\nicefrac{{1}}{{V}}\mathds{1}$triangleright Uniform reduction coefficient
7:     $\bm{\theta}_{v}\leftarrow\bm{\theta}^{(0)}$triangleright Standrad xLSTM initialization
@@
12:         $\bm{\phi}_{v}\leftarrow\bm{\phi}_{v}-\eta\bm{g}_{\bm{\phi}}$triangleright GD step
13:         $\bm{\theta}_{v}\leftarrow\bm{\theta}_{v}-\eta\bm{g}_{\bm{\theta}}$triangleright GD step
@@
15:              $\bm{\alpha}_{v}\leftarrow\bm{\alpha}_{v}-\eta\bm{g}_{\bm{\alpha}}$triangleright GD step
@@
17:         $\bm{W}_{v}\leftarrow\left(1-\frac{\lambda\eta\bm{\alpha}_{v}}{\left\lVert\bm{W}_{v}\right\rVert_{2}}\right)_{+}\bm{W}_{v}$triangleright Compression step with proximal GD
@@
21:$\mathcal{E}\leftarrow\left\{(v,w)\in\mathcal{V}\times\mathcal{V}\;\middle|\;\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}>0\right\}$
@@ Appendix B On the Approximation Capabilities of sLSTM blocks
Classic RNNs have long been known to be extremely powerful models of computation. Specifically, they are Turing complete [Siegelmann and Sontag, 1995] and, by extension, universal function approximators [Song et al., 2023]. Traditional LSTMs as proposed by Hochreiter and Schmidhuber [1997] are universal function approximators being at least as powerful as the RNNs [Song et al., 2023, Corollary 16]. Due to their novelty, the sLSTM cells as presented in Section 2.3 have not yet received the same degree of theoretical analysis. Yet, it appears natural to extend the same reduction to RNNs as has been shown for LSTMs, since the main technicality is the differing normalization of the hidden state [Beck et al., 2024, Eq. 36] and the alternative exponential activation function, which Song et al. [2023] abstract away with. sLSTM blocks then wrap these cells with additional operations (cf. Beck et al. [2024], App. A.4). They, however, can be carefully configured to reduce to only rescaling Layer Normalizations and residual connections, which are known to, again, yield universal function approximators for many architectures [Tabuada and Gharesifard, 2020, Marchi et al., 2021]. This can be shown by first omitting the anyway optional upfront convolution and Swish activation. We, furthermore, set the number of heads to $N_{h}=1$, causing the per-head Group Normalization after the cell to degenerate to yet another Layer Normalization. Lastly, the final post up-projection of $\bm{x}_{\text{in}}$ to $\bm{x}_{\text{out}}$ is defined as
@@ Appendix C Dataset Details
This section details all six datasets used in our empirical evaluation of GC-xLSTM.
@@
The $V$-dimensional Lorenz-96 model [Karimi and Paul, 2010] is a chaotic multivariate dynamical system governed by the differential equations $\frac{dx_{t,i}}{dt}=(x_{t,i+1}-x_{t,i-2})x_{t,i-1}-x_{t,i}+F,$ with the external forcing coefficient $F$ regulating the nonlinearity of the system. Low values of $F$ correspond to near-linear dynamics, while higher values induce chaotic behavior. There are two sources of randomness in the system. Firstly, we sample i.i.d. starting conditions from $\mathcal{N}(0,0.01)$. Secondly, in each step of the simulation, we add i.i.d. noise sampled from $\mathcal{N}(0,0.1)$ to $x_{t,i}$. Following the setting of Tank et al. [2022] for best comparability, we simulate $V=20$ variates with a sampling rate of $\Delta t=0.05$ for a total of $T=500$ time steps. We use two forcing constants $F\in\{10,40\}$ to test our model under different levels of non-linearity.
@@
Discovering connectivity networks within (human) brains is a key application of GC detection methods. To this end, brain activity is measured non-invasively using functional magnetic resonance imaging (fMRI) over time and grouped into regions between which connections are to be found. Specifically, we used the realistic simulations of blood-oxygen-level-dependent (BOLD) deconvolved data of Smith et al. [2011]. The data contains $V=15$ variates over $T=200$ time steps.
Moléne Dataset.
The Moléne dataset [Girault, 2015] contains hourly temperatures recorded by sensors at $V=32$ locations in Brittany, France, during $T=744$ hours. The objective is to understand the spatio-temporal dynamics of the temperature and to assess the extent to which the model can uncover complex relationships in weather by considering only local observations and their geographical positioning.
@@
We also apply our methodology to detect complex, nonlinear dependencies in human motion capture (MoCap) recordings. In contrast to the Lorenz-96 simulated dataset results, this analysis allows us to more easily visualize and interpret the learned network. We consider a data set from the CMU MoCap database [CMU, 2009]. The data recorded from 30 unique regions in the body consists of $V=54$ joint angle and body position recordings across multiple different subjects. Since some regions, like the neck, have multiple degrees of freedom in both translation and rotation, we consider the GC relations between two joints based on edges between all movement directions. The motion ranges from locomotion (e.g., running and walking) over physical activities such as gymnastics and dancing to day-to-day social interactions.
@@
Another common field of application of GC detection methods is discovering links in economic data, where it is otherwise hard to maintain an overview. We, therefore, benchmark GC-xLSTM on company fundamentals data [Divo et al., 2024]. The dataset contains $V=20$ economic variables such as the Net Income and the Total Liabilities of 2527 companies. The data was collected quarterly from 2009 Q1 to 2023 Q3, resulting in only $T=56$ time steps.
@@
Following the well-known setup of Tank et al. [2022], we generate a two-step linear auto-regressive process in $V\in\{10,20\}$ variates. All variables depend on themselves, and an additional three random dependencies are added as targets to be discovered. The number of time steps varies in $T\in\{250,500,1000\}$.

### v5

@@ 5 Conclusion
This work received funding from the ACATIS Investment KVG mbH project “Temporal Machine Learning for Long-Term Value Investing” and the KompAKI project of the German Federal Ministry of Research, Technology and Space within the “The Future of Value Creation – Research on Production, Services and Work” program (funding number 02L19C150), managed by the Project Management Agency Karlsruhe (PTKA). The TU Eindhoven author received support from their Dep. of Mathematics and Computer Science and the Eindhoven AI Systems Institute. Furthermore, this work benefited from the DYNAMIC Centre funded by the LOEWE program of the Hessian Ministry of Science and Research, Arts and Culture (HMWK) as LOEWE 1/16/519/03/09.001 (0009)98 and the HMWK project “The Third Wave of Artificial Intelligence – 3AI”. It also benefited from the early stage of the German Federal Ministry for Economic Affairs and Energy project “Souveräne KI für Europa” (13IPC040G) as part of the EU funding program IPCEI-CIS; funding has not started yet, and from early stages of the Cluster of Excellence “Reasonable AI” funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy – EXC-3057; funding will begin in 2026. Funding for H. Poonia to attend NeurIPS was provided by the CMU GSA/Provost Conference Funding. The authors are responsible for the content of this publication. Map data © OpenStreetMap contributors, licensed under the ODbL and available from openstreetmap.org.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We explicitly list the contributions at the end of Section˜1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations as an explicitly named paragraph in Section˜5. We acknowledge that any empirical results are by their very nature limited to the settings in which they were obtained, and thus strive to accurately describe them for best reproducibility (see also Question 4).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: While we perform theoretical anlaysis in Sections˜3.3 and B, these are not in the form of mathematical theorems.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All implementation details, including dataset descriptions and experiment configurations, are provided in Section˜4.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We strive to use freely available datasets where possible and exclusively employ openly available software. Furthermore, we provide the source code for full reproducibility at github.com/harpoonix/GC-xLSTM.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All such details are provided in Section˜4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All quantitative results are accompanied by standard deviations (specifically, see Table˜1).
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: This is provided at the beginning of Section˜4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We made sure to comply with the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: While we acknowledge that many technologies have wide-ranging societal impacts, our primary focus is on technical innovation. We have thus not identified specific concerns requiring emphasis in this work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Please see Question 10. Specifically, we do not provide any trained models or similar high-risk artifacts.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We carefully cite all immediately relevant scholarly works and provide URLs to any other resources in Sections˜4 and C.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The documentation of the source code is also provided at github.com/harpoonix/GC-xLSTM.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We did not perform any such experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Please see Question 14.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We did not employ LLMs in any part of GC-xLSTM.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
@@ Appendix A Full Algorithm
Algorithm˜1 describes the entire procedure for extracting a Granger causal graph using GC-xLSTM. $(\cdot)_{+}$ denotes truncation as $\max(\cdot,0)$.
@@
5:  $\bm{\phi}_{v}:\bm{W}_{v},\bm{b}_{v}\sim\mathcal{U}\left(-\sqrt{\nicefrac{{1}}{{V}}},\sqrt{\nicefrac{{1}}{{V}}}\right)$ $\triangleright$ Kaiming/He initialization
6:  $\bm{\alpha}_{v}\leftarrow\nicefrac{{1}}{{V}}\mathds{1}$ $\triangleright$ Uniform reduction coefficient
7:  $\bm{\theta}_{v}\leftarrow\bm{\theta}^{(0)}$ $\triangleright$ Standard xLSTM initialization
@@
12:   $\bm{\phi}_{v}\leftarrow\bm{\phi}_{v}-\eta\bm{g}_{\bm{\phi}}$ $\triangleright$ GD step
13:   $\bm{\theta}_{v}\leftarrow\bm{\theta}_{v}-\eta\bm{g}_{\bm{\theta}}$ $\triangleright$ GD step
@@
15:     $\bm{\alpha}_{v}\leftarrow\bm{\alpha}_{v}-\eta\bm{g}_{\bm{\alpha}}$ $\triangleright$ GD step
@@
17:   $\bm{W}_{v}\leftarrow\left(1-\frac{\lambda\eta\bm{\alpha}_{v}}{\left\lVert\bm{W}_{v}\right\rVert_{2}}\right)_{+}\bm{W}_{v}$ $\triangleright$ Compression step with proximal GD
@@
[…]
@@ Appendix B On the Approximation Capabilities of sLSTM blocks
Classic RNNs have long been known to be extremely powerful models of computation. Specifically, they are Turing complete [Siegelmann and Sontag, 1995] and, by extension, universal function approximators [Song et al., 2023]. Traditional LSTMs as proposed by Hochreiter and Schmidhuber [1997] are universal function approximators being at least as powerful as the RNNs [Song et al., 2023, Corollary 16]. Due to their novelty, the sLSTM cells as presented in Section˜2.3 have not yet received the same degree of theoretical analysis. [^2] A discussion of expressivity hierarchies can be found in Auer et al. [2025, App. A]. Yet, it appears natural to extend the same reduction to RNNs as has been shown for LSTMs, since the main technicality is the differing normalization of the hidden state [Beck et al., 2024, Eq. 36] and the alternative exponential activation function, which Song et al. [2023] abstract away with. sLSTM blocks then wrap these cells with additional operations (cf. Beck et al. [2024], App. A.4). They, however, can be carefully configured to reduce to only rescaling Layer Normalizations and residual connections, which are known to, again, yield universal function approximators for many architectures [Tabuada and Gharesifard, 2020, Marchi et al., 2021]. This can be shown by first omitting the optional upfront convolution and Swish activation. We, furthermore, set the number of heads to $N_{h}=1$, causing the per-head Group Normalization after the cell to degenerate to yet another Layer Normalization. Lastly, the final post up-projection of $\bm{x}_{\text{in}}$ to $\bm{x}_{\text{out}}$ is defined as
@@ Appendix C Dataset Details
This section details all six datasets used in our empirical evaluation of GC-xLSTM. An overview is provided in Table˜5.
Table 5: Overview of the diverse collection of datasets. It also shows the number of variates $V$, time steps $T$, samples NN, and look-back context steps of GC-xLSTM $C$.
Name Origin Type $V$ $T$ NN $C$
Lorenz-96 Karimi and Paul [2010] Simulated 20 500 1 10
fMRI Smith et al. [2011] Real-world 15 200 1 10
Moléne Girault [2015] Real-world 32 744 1 10
Human MoCap (Run) CMU [2009] Real-world 54 1232 61 10
Human MoCap (Salsa) CMU [2009] Real-world 54 4136 30 10
Company Fundamentals Divo et al. [2025] Real-world 19 56 2527 40
VAR Karimi and Paul [2010] Simulated 10/20 mult. 1 5
@@
The $V$-dimensional Lorenz-96 model [Karimi and Paul, 2010] is a chaotic multivariate dynamical system governed by the differential equations
$\frac{dx_{t,i}}{dt}=(x_{t,i+1}-x_{t,i-2})x_{t,i-1}-x_{t,i}+F,$
with the external forcing coefficient $F$ regulating the non-linearity of the system. Low values of $F$ correspond to near-linear dynamics, while higher values induce chaotic behavior. There are two sources of randomness in the system. Firstly, we sample i.i.d. starting conditions from $\mathcal{N}(0,0.01)$. Secondly, in each step of the simulation, we add i.i.d. noise sampled from $\mathcal{N}(0,0.1)$ to $x_{t,i}$. Following the setting of Tank et al. [2022] for best comparability, we simulate $V=20$ variates with a sampling rate of $\Delta t=0.05$ for a total of $T=500$ time steps after a brief burn-in time. We use two forcing constants $F\in\{10,40\}$ to test our model under different levels of non-linearity.
@@
Discovering connectivity networks within (human) brains is a key application of GC detection methods. To this end, brain activity is measured non-invasively using functional magnetic resonance imaging (fMRI) over time and grouped into regions between which connections are looked for. Specifically, we used the realistic simulations of blood-oxygen-level-dependent (BOLD) deconvolved data of Smith et al. [2011].
Moléne.
The Moléne dataset [Girault, 2015] contains hourly temperatures recorded by sensors at $V=32$ locations in Brittany, France, during $T=744$ hours. The objective is to understand the spatio-temporal dynamics of the temperature and to assess the extent to which the model can uncover complex relationships in weather by considering only local observations.
@@
We also apply our methodology to detect complex, nonlinear dependencies in human motion capture (MoCap) recordings. In contrast to the Lorenz-96 simulated dataset results, this analysis allows us to visualize and interpret the learned network more easily. We consider a data set from the CMU MoCap database [CMU, 2009]. The data comprises $V=54$ joint angle and body position recordings across multiple subjects. Since some regions, like the neck, have multiple degrees of freedom in both translation and rotation, we consider the GC relations between two joints based on edges between all movement directions. The motion ranges from locomotion (e.g., walking) over physical activities such as gymnastics and dancing to day-to-day social interactions.
@@
Another common field of application of GC detection methods is discovering links in economic data, where it is otherwise hard to maintain an overview. We, therefore, benchmark GC-xLSTM on company fundamentals data [Divo et al., 2025]. The dataset contains $V=19$ economic variables, such as the Net Income and the Total Liabilities of 2527 companies. The data was collected quarterly from 2009 Q1 to 2023 Q3, resulting in only $T=56$ time steps.
@@
Following the well-known setup of Tank et al. [2022], we generate a two-step linear auto-regressive process in $V\in\{10,20\}$ variates. All variables depend on themselves, and an additional three random dependencies are added as targets to be discovered. The number of time steps varies in $T\in\{250,500,1000\}$, and we start recording the data after a brief burn-in time.